OpenStudy (anonymous):

If f(x) = 5x – 3 and g(x) = 3x – 3, find f(x) – g(x). A. 2x B. 8x – 6 C. 2x – 6 D. 8x

OpenStudy (anonymous):

@yrelhan4 Would You Mind Explaining This One ?

OpenStudy (yrelhan4):

Just subtract the two expressions. Simple.

OpenStudy (anonymous):

How Do I Go By Doing That ?

OpenStudy (anonymous):

2x-6?

OpenStudy (yrelhan4):

f(x) = 5x – 3 and g(x) = 3x – 3 The question asks you for f(x) - g(x).. So, (5x-3) - (3x-3) Just solve that.

OpenStudy (yrelhan4):

No. 5x - 3 - 3x + 3.. So you get 2x.

OpenStudy (anonymous):

f(x)=5x-3_g(x)=3x-3_f(x)-g(x) f(x(x)=5x^(2)-15 (5x^(2)-15) 5x^(2)-15

OpenStudy (anonymous):

The Question Is If f(x) = 5x – 3 and g(x) = 3x (– 3), find f(x) – g(x). They Are Both -3?

OpenStudy (yrelhan4):

Eh, wrong post @Luis_Rivera ?

OpenStudy (yrelhan4):

But you also have a minus sign outside the bracket. So minus(outside the bracket) with minus(with 3 inside the bracket) makes it a plus.

OpenStudy (anonymous):

But The -3 Is On The Outside?

OpenStudy (yrelhan4):

-3 is inside the expression of g(x). so when the question asks you for [f(x)]-[g(x)].. it asks you for (5x – 3) - (3x-3)